# dMel: Speech Tokenization made Simple

### v1

@@ 5 Conclusion
In this work, we proposed dMel, a novel train-free speech tokenization method that discretizes mel-filterbank energies directly into bins. By operating on the authentic mel-filterbank representation, dMel inherently preserves both semantic and acoustic information in a unified tokenized representation. Our key contribution is the evaluation of dMel within a unified transformer decoder-only architecture for speech recognition (ASR) and speech synthesis (TTS) tasks. Our dMel-based ASR model, RichASR, achieved the lowest word error rate among tokenization methods, robustly preserving semantic content. For TTS, dMel’s generation yielded the lowest WER, accurately reconstructing speech waveforms. Our dMel-based TTS model, RichTTS, achieved competitive naturalness, lowest error rates, and long audio generation capabilities.
dMel’s simplicity circumvents separate tokenizers or multi-stage modeling, reducing computational overhead and dependence on pretrained models. By unifying semantic and acoustic modeling, dMel enables efficient speech-text modeling frameworks. While initial joint TTS-ASR training showed promise, further work is needed. Our primary contribution demonstrates dMel’s effectiveness for high-performing separate TTS and ASR models within a unified decoder architecture.
6 Limitations
Because TTS work is tremendously fragmented and clear protocols are not often available for training and evaluation, we reimplemented other tokenizers within our code base using publicly available, official implementations where available. We trained our models on those tokenizations. While we made the best effort to tune the tokenization methods and the models, there is always a possibility we missed some details. However, our results seem to tell a consistent story when viewed from multiple angles, and when viewed on multiple datasets. For the joint model, we did not do extensive multi-task training using text only data and we intend to do that in the future work. We also did not train on larger model sizes (>1B parameters), larger datasets (>1k hours), or using pretrained models.
7 Acknowledgements
We thank Dan Busbridge, Ronan Collobert and Jason Ramapuram for their helpful feedback and critical discussions at all stages of the research; Rick Chang, David Grangier, Barry Theobald for their helpful feedback on the initial paper draft. Names are in alphabetical order by last name within group.
@@ Appendix A Ethics Statement
Appendix B Data, Code, Reproducibility
@@
Appendix C Subjective Evaluation for TTS
@@
Appendix D Training Details
For our decoder-only model we stack together speaker embedding, speech tokens and text tokens. Both speech and text tokens have prepended begin of sentence token (<bos>) and appended end of sentence token (<eos>).
We train all models using the Adam optimizer with a learning rate of 1e-3, learning rate warmup of 4k steps for ASR and 5k for TTS, cosine learning rate schedule and gradient clipping of 1.0 for TTS and 0.1 for ASR and joint models. We use dynamic batching to optimize the data packing with total batch size of 1.4h/1.4h/0.7h for ASR training and 1h/2h/2h for TTS training for Small/Base/Large models. We train TTS models for 600k steps and ASR models 80k steps with mixed precision training and BF16 on A100 and H100 GPUs with 80GB. Both ASR models and TTS models are trained with 8GPUs for less than a day and for 2-4 days for ASR and TTS respectively.
Table 9: Decoder-only transformer model configurations for ASR, TTS and joint models training.
@@ Appendix D Training Details
D.1 LM-Style Speech-to-Text
@@
During experiments with ASR decoder-only models and dMel tokenization we observed some model training instabilities resulting in the spikes in the gradient norm and thus sometimes spikes in training loss. To stabilize training i) we reduce the gradient clipping from 1.0 used for TTS training to 0.1; ii) we add queries and keys normalizations via LayerNorm [12] on the head dimension before computing attention matrix.
@@
As we found one best model configuration for the Base model with dMel we then change only i) model size ii) speech tokenization iii) training data (here we increase model dropout to 0.3 for training on train-clean-360 and to 0.5 for training on train-clean-100 as otherwise models drastically overfits.); the rest of hyper-parameters stays the same.
Appendix E Ablations
E.1 LM-Style Text-to-Speech
@@
E.2 LM-Style Speech-to-Text
Table 11: Our ASR models trained on different subsets (train-clean-100 LS-100, train-clean-360 LS-360, full LibriSpeech LS-960) of LibriSpeech, with different model size and different speech tokenization (greedy decoding is reported). Results are shown across 2 runs with mean WER and standard deviation.
tokenization model size data dev-clean dev-other test-clean test-other

### v2

@@ 5 Conclusion
In this work, we proposed dMel, a novel train-free speech tokenization method that discretizes log mel-filterbank energies directly into bins. By operating on the authentic log mel-filterbank representation, dMel inherently preserves both semantic and acoustic information in a unified tokenized representation. Our key contribution is the evaluation of dMel within a unified LM-style transformer architecture for speech recognition (ASR) and speech synthesis (TTS) tasks. Our dMel-based ASR model, RichASR, achieved the lowest word error rate among tokenization methods, robustly preserving semantic content. For TTS, dMel’s generation yielded the lowest WER, accurately reconstructing speech waveforms. Our dMel-based TTS model, RichTTS, achieved competitive naturalness, lowest error rates, and long audio generation capabilities.
dMel’s simplicity circumvents separate tokenizers or multi-stage modeling, reducing computational overhead and dependence on pretrained models. By unifying semantic and acoustic modeling, dMel enables efficient speech-text modeling frameworks. While initial joint TTS-ASR training showed promise, further work is needed. Our primary contribution demonstrates dMel’s effectiveness for high-performing separate TTS and ASR models within a unified LM-style architecture.
@@ Appendix A Ethics Statement
Appendix B Limitations
Because TTS work is tremendously fragmented and clear protocols are not often available for training and evaluation, we reimplemented other tokenizers within our code base using publicly available, official implementations where available: e.g. we used Hubert-KM and speech tokenizer features extraction from the public codebases and pluged them into our LM-style model training. While we made the best effort to tune the tokenization methods and the models, there is always a possibility we missed some details. However, our results seem to tell a consistent story when viewed from multiple angles, and when viewed on multiple datasets. We also did not train on larger model sizes (>1B parameters), larger datasets (>1k hours), or using pretrained models.
The real challenge for modern multimodal LLMs is complex semantic understanding tasks. While our current experiments focus on text-to-speech and speech-to-text tasks, these encompass critical aspects of speech processing. dMel’s effective performance within a decoder-only architecture for both tasks suggests potential for broader applications. We recognize the importance of more sophisticated speech understanding tasks and view our work as a foundation for future research leaving other tasks out of scope of the paper. Scaling up pretraining and exploring complex semantic understanding tasks could further validate our approach’s versatility across a wider range of multimodal language processing challenges.
We acknowledge that our current scope targets only speech on purpose, as indicated in our title. While dMel may potentially support non-speech tasks, our current exploration and verification focus solely on speech, not general audio. Regarding the “speaker variations” – mel-spectrogram is used for speaker recognition widely, thus it preserves necessary speaker information on which we thus rely in dMel too.
Appendix C Data, Code, Reproducibility
@@
Appendix D Subjective Evaluation for TTS
@@
We noticed human annotators have bias over audio volume so we do volume normalization on top of all reconstructed or generated audio before giving them to human annotators.
We report Mean Opinion Score (MOS) results throughout the paper with confidence intervals calculated using bootstrap resampling with 1000 iterations, providing a reliable estimate of the variability MOS results.
Appendix E Training Details
E.1 Baselines
E.2 RichASR and RichTTS
For our LM-style model we stack together speaker embedding, speech tokens and text tokens. Both speech and text tokens have prepended begin of sentence token (<bos>) and appended end of sentence token (<eos>).
We train all models using the Adam optimizer with a learning rate of 1e-3, learning rate warmup of 4k steps for ASR and 5k for TTS, cosine learning rate schedule and gradient clipping of 1.0 for TTS and 0.1 for ASR and joint models. We use dynamic batching to optimize the data packing with total batch size of 1.4h/1.4h/0.7h for ASR training and 1h/2h/2h for TTS training for Small/Base/Large models. We train TTS models for 100k steps and ASR models 80k steps with mixed precision training and BF16 on A100 and H100 GPUs with 80GB. Both ASR models and TTS models are trained with 8GPUs for less than a day and for 2-4 days for ASR and TTS respectively.
Table 12: LM-style transformer model configurations for ASR, TTS and joint models training.
@@ E.2 RichASR and RichTTS
E.3 LM-Style Speech-to-Text
@@
As we found one best model configuration for the Base model with dMel we then change only i) model size ii) speech tokenization iii) training data (here we increase model dropout to 0.3 for training on train-clean-360 and to 0.5 for training on train-clean-100 as otherwise models drastically overfit); the rest of hyper-parameters stay the same.
Appendix F Ablations
F.1 LM-Style Text-to-Speech
@@
F.2 LM-Style Speech-to-Text
ASR ablations for different model sizes, data sizes, and tokenizers are shown in Table 14.
We noticd the results in [10] seems to be the SOTA for LM-style ASR model to the best of our knowledge. However, as many ablations are missed in [10], we took their open-sourced code and run ablations ourselves to have proper comparison with it. The final results, including ablation with dMel are shown in Table 15:
  • We successfully reproduced [10] results (row 1 and 2).
  • Without pretraining (rows 3, 4, 5):
    dMel outperforms HuBERT-KM on both clean and other datasets; dMel surpasses BPE on top of HuBERT-KM on clean data, while BPE on HuBERT-KM performs better on other.
  • Without pretraining and without speed perturbation (rows 6, 7, 8):
    BPE on HuBERT-KM performance decreases significantly after diabling speed perturbation (compare rows 3 and 6), raising questions about its generalizability to other domains, given that BPE tokens are trained on speed-perturbed LibriSpeech data.
    Our dMel (row 8) achieves substantially better results than both HuBERT-KM and BPE on HuBERT-KM (rows 7 and 6), demonstrating robust performance even without speed augmentation.
Note that in dMel, we use SpecAugment (masking across time and channels) and [10] also use SpecAugment. According to their code, the time masking is 30%, while channel masking is impossible as there is only 1 channel).
We believe these results demonstrate the effectiveness, simplicity in use, and robustness of our dMel tokenization method, particularly in scenarios where extensive pretraining or domain-specific augmentations may not be feasible.
Note that [10] did not show applicability of BPE on HuBERT-KM or HuBERT-KM to TTS task, while in VOXTLM (also uses BPE on HuBERT-KM) it is shown that this tokenization is not suited for TTS (the performance is poor). dMel in contrary is shown to perform well on TTS task too in addition to ASR.
Table 14: Our ASR models trained on different subsets (train-clean-100 LS-100, train-clean-360 LS-360, full LibriSpeech LS-960) of LibriSpeech, with different model sizes and different speech tokenizations (greedy decoding is reported). Results are shown across 2 runs with mean WER and standard deviation.
Tokenization Model Size Data dev-clean↓ dev-other↓ test-clean↓ test-other↓
@@ F.2 LM-Style Speech-to-Text
Table 15: Ablations on the LM-style ASR model with GPT-2 architecture using the setup from [10]: we ablate pretraining with text, speed perturbation and speech tokenization methods. All models are trained on LibriSpeech 960h. We report WER (%) on LibriSpeech validation and test sets.
Tokenization Text Pretraining Speed Perturbation dev-clean↓ dev-other↓ test-clean↓ test-other↓
BPE on HuBERT+KM, [10] 2.9 6.2 3.0 6.6
BPE on HuBERT+KM, reproduction 2.9 6.3 3.2 6.7
BPE on HuBERT+KM 3.6 7.8 3.8 8.3
HuBERT+KM 5.1 8.9 5.5 9.3
dMel 3.1 8.4 3.4 8.6
BPE on HuBERT+KM 5.4 9.7 5.3 10.0
HuBERT+KM 6.0 10.5 6.4 10.9
dMel 3.7 9.7 3.9 9.5
F.3 Joint Speech-Text Modeling Discussion
We found it to be challenging to train joint model for ASR and TTS, similar to observations as in [24] and e.g. [34, 31] for joint audio-visual speech recognition. Also, there is a very recent research work [36], that also shows training TTS and ASR jointly is challenging, and needs carefully designed model architecture and training loss fusion technique.
One of the reasons is the different pace of learning. Careful consideration of training strategies can mitigate some of the challenges in joint modeling of TTS and ASR tasks, highlighting the complexities inherent in combining these distinct but related tasks within a single model.
Another reason we suspect is the mismatch between train and test time, which is more pronounced for the joint modeling: if we compare individual validation losses per task in joint model to their one-task training counterparts we see they match each other (so training is fine), however the generation (test time which mismatches how the train loss is defined) for both tasks is broken: longer sequences has hallucination and high repetition issues. This could be due to different length of sequences between text and audio and thus learnt attention pattern could be different which creates longer sequences generation issue for the joint model.
Last but not the least, the two tasks have opposite modalities in the input and output, making it rather difficult to model. Most previously researched multi-task work have the same modality in the output. The combination of ASR and TTS is a rather recent phenomenon, such as Viola and VOXTLM.